[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Nov 10, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Nov 10 18:35:09 UTC 2016


HSU Shift Log for Thu Nov 10 2016 13:35:02 GMT-0500 (EST)

Shift: H

Forecaster: R Berg

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues: Both the 06Z and 12Z cycles of the GFS were slow  
coming in. Craig indicated that we've been placed on an older router while  
they do an upgrade, and the runs are about 1-1.25 hrs late. This should go  
back to normal once the upgrade is complete.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161110/3619c1bf/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list